OpenStudy (anonymous):

Solve using the substitution method. Show your work. If the system has no solution or an infinite number of solutions, state this. 4a + 20b = 8 8a = -40b + 16

OpenStudy (anonymous):

This system has no solution: as ... ... the two equations are identical ... 4a + 20b = 8 (1) 8a = -40b + 16 (2) divide (2) by 2 and rearange ... u get 4a+20b =8 which ='s (1)

OpenStudy (anonymous):

4a = 8 - 20b 8a = 16 - 40b + ----------------- 12a = 24 - 60b .... ...

OpenStudy (anonymous):

12a + 60b = 24 :D

OpenStudy (anonymous):

and yea this equation systems solution is not foundable, 4a = 8 - 20b = 8a = 16 - 40b so we can't find the result :D there should be a differend eqution to find the result :D

OpenStudy (anonymous):

there is no solution for this system of equation because the never have an intersection poin as they represt 2 parallel lines.